OpenStudy (anonymous):

If f(x) = x minus 2, all over 2, then what is f–1(x) ?

OpenStudy (johnweldon1993):

\[y = \frac{ x - 2 }{ 2 }\] F(x)^-1 means the inverse....here you take your original equation (above) and switch the 'x' and the 'y' ...so it becomes \[x = \frac{ y - 2 }{ 2 }\] Now solve for 'y' again...can you do that...?

OpenStudy (anonymous):

No, i'm so beyond confused

OpenStudy (johnweldon1993):

Okay don't worry about it....so we have \[x = \frac{ y - 2 }{ 2 }\] We want to solve for 'y' again... so first....we want to get rid of that annoying fraction...since everything is being divided by 2....we should cancel that out...and multiply by 2 2x = y - 2 Notice how everything divided by 2....when multiplied by 2...cancels out... so now that all you have is 2x = y - 2 what do you need to do to solve for 'y'...?

OpenStudy (anonymous):

add 2 too both sides

OpenStudy (johnweldon1993):

Correct...so out final answer would be y = 2x + 2 That is the inverse

OpenStudy (anonymous):

thank you!

OpenStudy (johnweldon1993):

No problem!